OpenStudy (anonymous):

Solve the equation 1/5x + 2 = 8 for x. 50 30 40 10

OpenStudy (turingtest):

1)subtract 2 from both sides 2)multiply both sides by 5

OpenStudy (anonymous):

you can multiply through by 5x so you have: 1+10x = 40x and then solve for x: x = 1/30

OpenStudy (anonymous):

THANK YOU I took forever to figure it out how come theres sooo many genius people on this site and Im not one of them:(

OpenStudy (anonymous):

whay how estebananya

OpenStudy (anonymous):

what*

OpenStudy (anonymous):

He is trying to show you how to clear the fraction

OpenStudy (anonymous):

how do you do that

OpenStudy (anonymous):

multiply both sides by 5 in this case, if their were more than one fraction you would have to find a comman denomantor.

OpenStudy (turingtest):

More importantly, why? You turn a two-step problem into a three-step one that way...

OpenStudy (anonymous):

you could also just divide 6 by 1/5th and get the same answer

OpenStudy (anonymous):

I did that I end up having 1x+2=40 thenwhat

OpenStudy (turingtest):

what do you get just subtracting 2 from both sides? 1/5x+2-2=8-2 simplify above...

OpenStudy (anonymous):

you would end up with x+10=40, x=30...but it's not important test had the quickest way. Just be sure you know the concepts behind what your doing

OpenStudy (anonymous):

where does the ten come in

OpenStudy (anonymous):

is it (1/5)x or 1/(5x)? if the answer is the first, i messed up

OpenStudy (anonymous):

can some one just show me step by step NOT missing any steps how to do it cuz I REALLY want to learn....

OpenStudy (anonymous):

before you worry about clearing functions or anything like that make sure you get your x on one side and your constant on the other.

OpenStudy (anonymous):

then you can multiply by 5 inorder to clear your function and solve the expresion in this case.

OpenStudy (anonymous):

Or if you don't feel like clearing a function divide by the coificent infront of the varible, which i do most of the time unless i don't have a calculator.

OpenStudy (turingtest):

1)subtract 2 from both sides: 1/5x+2=8 -2 -2 1/5x=6 2) multiply both sides by 5. I cannot do this for you as it directly gives the answer.

OpenStudy (anonymous):

how are you'll sooo smart:( thankx

OpenStudy (anonymous):

Lots of practice

OpenStudy (anonymous):

I still can't help but feel like there's some ambiguity. Is it \[\frac{1}{5x} + 2 = 8\] or \[\frac{1}{5}x + 2 = 8\]?

OpenStudy (turingtest):

Given the answer choices there is no ambiguity though, eh?

OpenStudy (anonymous):

the second one Jemurray3

OpenStudy (anonymous):

Oh, I didn't even read the choices... nevermind me :) In the future baracuda, just in case, use parentheses if you can.

OpenStudy (anonymous):

ok